Consider an asset that costs $730,000 and is depreciated straight-line to zero over its seven-year tax life. The asset is to be used in a five-year project; at the end of the project, the asset can be sold for $163,000. If the relevant tax rate is 24 percent, what is the aftertax cash flow from the sale of this asset? (Do not round intermediate calculations.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useul Life

=(730,000/7)=$104,285.7143/year

Hence book value as on date of sale=Cost-Accumulated depreciation

=730,000-(104,285.7143*5)

=$208,571.4285

Hence since sale proceeds is less than book value;loss on sale=($208,571.4285-$163000)

=$45571.4285

Hence aftertax cash flow=Sale proceeds+(loss on sale*Tax rate)

=163000+(45571.4285*0.24)

=$173937.14(Approx).